Determine the volume of SO2 (at STP) formed from the reaction of 96.7 g of FeS2 and 55.0 L of O2 (at 398 K and 1.20 bar). The molar mass of FeS2 is 119.99 g mol-1.

4FeS2(s) + 11O2(g) → 2Fe2O3(s) + 8SO2(g)
1
Expert's answer
2020-12-23T04:07:01-0500

R = 8.314×10−2 L⋅bar⋅K−1⋅mol−1 (universal gas constant).


Solution:

The balanced chemical equation:

4FeS2(s) + 11O2(g) → 2Fe2O3(s) + 8SO2(g)

According to the equation: n(FeS2)/4 = n(O2)/11 = n(SO2)/8


1) Moles of FeS2 = Mass of FeS2 / Molar mass of FeS2

The molar mass of FeS2 is 119.99 g mol-1.

Hence,

n(FeS2) = Moles of FeS2 = 96.7 g / 119.99 g mol-1 = 0.806 mol

n(FeS2) = 0.806 mol

2) According to the ideal gas law: PV = nRT

Moles of O2 = PV / RT

n(O2) = Moles of O2 = (1.20 bar × 55.0 L) / (8.314×10−2 L⋅bar⋅K−1⋅mol−1 × 398 K ) = 1.9946 mol

n(O2) = 1.9946 mol


According to stoichiometry:

11 moles of O2 reacts with 4 moles of FeS2

Thus 1.9946 moles of oxygen reacts with:

(1.9946 mol O2 × 4 mol FeS2) / 11 mol O2 = 0.7253 mol FeS2.

However, initially there is 0.806 mol of FeS2 (according to the task).

Thus O2 acts as limiting reagent and FeS2 is excess reagent.


Therefore,

n(O2)/11 = n(SO2)/8

n(SO2) = 8 × n(O2) / 11 = (8 × 1.9946 mol) / 11 = 1.4506 mol SO2

n(SO2) = 1.4506 mol


At STP, one mole of any gas occupies a volume of 22.4 L.

Thus 1.4506 moles of SO2 occupies:

(1.4506 mol × 22.4 L) / 1 mol = 32.49 L of SO2

V(SO2) = 32.49 L = 32.5 L (at STP).


Answer: 32.5 L of SO2 are formed.